# Rethinking Attention: Polynomial Alternatives to Softmax in Transformers

### v2

@@ 3 Preliminaries and Notation
The factor $\frac{1}{\sqrt{d}}$, as explained in [26], is a scaling to prevent the gradients of softmax from being too small. For the theoretical analysis in this paper we will only use the dot-product similarity $qk^{T}$ and call the $N\times N$ matrix $softmax(qk^{T})$ the softmax self-attention matrix. In the experiments, Section 6, we will empirically validate our theoretical framework on more general softmax attention blocks used in state of the art transformers such as DeiT [25], Swin Transformer [17] and XciT [30].
@@
Given a matrix $A\in\mathbb{R}^{n\times m}$, we denote its Frobenius norm by $||A||_{F}$. Additionally, we use the notation $\mathbb{E}$ to represent the expectation of a random variable, where the specific random variable being considered will be clear from the context.
@@ 4.1 Implicit regulatization of Softmax
$||\mathbf{softmax}(A)||_{F}$ $\leq\sqrt{N}$ (6)
$||\mathbf{\nabla softmax}(A)||_{F}$ $\leq 2\sqrt{N}.$ (7)
The key implication of theorem 4.1 is that during the training of a transformer with softmax self-attention, the Frobenius norm of each softmax self-attention matrix remains bounded by a value that grows as $\mathcal{O}(\sqrt{N})$. This ensures that backpropagation through the weights of the self-attention matrix does not lead to excessively large gradients. The proof hinges on the fact that the row normalization inherent in softmax effectively controls the Frobenius norm. For a detailed proof see Section A.1.1.
@@ 4.2 Polynomial activations for self-attention
In Section 4.1, we demonstrated that softmax implicitly regularizes the Frobenius norm of the self-attention matrix. Building on this, we now show that by scaling specific polynomial activations, a similar regularization effect on the Frobenius norm can be achieved in expectation, closely replicating the impact of softmax.
@@
$\mathbb{E}\bigg{|}\bigg{|}\bigg{(}\frac{XQK^{T}X^{T}}{\sqrt{d}}\bigg{)}^{p}\bigg{|}\bigg{|}_{F}\leq\mathcal{O}(N)$ (8)
@@
$\mathbb{E}\bigg{|}\bigg{|}\frac{1}{\sqrt{N}}\bigg{(}\frac{XQK^{T}X^{T}}{\sqrt{d}}\bigg{)}^{p}\bigg{|}\bigg{|}_{F}\leq\mathcal{O}(\sqrt{N}).$ (9)
Corollary 4.3 establishes that activations of the form $\phi(x):=\frac{1}{\sqrt{N}}x^{p}$ provide a level of regularization, in expectation, similar to that of softmax when applied to the self-attention matrix. The proof of theorem 4.2 can be found in appendix A.1.2. The next property we want to prove is one similar to the gradient bound obtained in Theorem 4.1. Since the self-attention matrix has parameters given by the queries $Q$ and keys $K$ [26], this implies that during the training of a transformer the $Q$ and $K$ matrices are the only aspects of the self-attention matrix that get updated. Therefore, we compute a derivative bound with respect to the $Q$ and $K$ derivatives.
@@
$\mathbb{E}\bigg{|}\bigg{|}\frac{\partial}{\partial Q}\bigg{(}\frac{(XQK^{T}X^{T})^{p}}{\sqrt{d}}\bigg{)}\bigg{|}\bigg{|}\leq\mathcal{O}(N)$ (10)
@@
$\mathbb{E}\bigg{|}\bigg{|}\frac{1}{\sqrt{N}}\frac{\partial}{\partial Q}\bigg{(}\frac{(XQK^{T}X^{T})^{p}}{\sqrt{d}}\bigg{)}\bigg{|}\bigg{|}\leq\mathcal{O}(\sqrt{N}).$ (11)
@@
Corollary 4.3 and Corollary 4.5 suggest that polynomial activations of the form $\phi(x)=\frac{1}{\sqrt{N}}x^{p}$, with $p>0$, can achieve performance comparable to softmax when applied to self-attention matrices. We point out that both corollaries rested on the assumption that $X$, $Q$ and $K$ are i.i.d random variables. In general, this is only true at initialization when training a transformer [2]. Although this is a limitation in the theory the experiments in Section 6 show that this insight can be used effectively to develop new attention blocks that perform comparable to softmax yet violate the three conditions of positivity, normalized rows summing to $1$ and sparsity showing that attention blocks do not need to be modeled as a probability distribution.
@@ 5 Testing the theory
In this section we test the theory developed in Section 4.2 on small vision transformers. We will consider the activation $\phi(x)=\frac{1}{k}x^{3}$, where $k>0$ is a fixed scale, as this activation clearly violates the three key conditions of softmax based attention; positivity, normalization and sparsity. We found that polynomials $\frac{1}{k}x^{p}$ for $p>3$ did not perform well during training as they witnessed a gradient vanishing problem due to the fact that the function $x^{p}$ for $p$ large have very small values around 00.
@@
The first experiment we conducted was to test how the Top-$1\%$ accuracy changes for a ViT-Tiny vision transformer [22], trained on Tiny-Imagenet dataset [21], as we change the sequence length $N$ of the input and the scale predicted in corollaries 4.3 and 4.5 when using the activation $\frac{1}{k}x^{3}$. The standard ViT-Tiny model comprises $12$ transformer layers, each equipped with $3$ attention heads, with each head having dimension $64$. We considered four different input sequence lengths $N$ of sizes $256$, $64$, $16$ and $8$. For each such sequence length, we ran a ViT-Tiny architecture with the activation $\frac{1}{k}x^{3}$ where $k$ ranged from roughly $10^{-3}$ to $10^{3}$. According to the theory developed in section 4, the Frobenius norm of $\frac{1}{\sqrt{N}}x^{3}$ scales according to $\mathcal{O}(\sqrt{N})$. Thus the best accuracy should occur when $k=\mathcal{O}(\sqrt{N})$ and should degrade for other values due to training instability. Figure 1 shows the results of the experiment, we note that the $x-axis$ plots the values of $k$. We see that as the sequence length increases the factor of $k$ needs to increase so that the activation $\frac{1}{k}x^{3}$ performs well on the ViT-Tiny architecture as predicted by the theory developed in corollaries 4.3 and 4.5.
@@
In a second experiment we decided to compare the performance of the original ViT-Tiny architecture, that uses an input sequence length of $256$, with a softmax activation and a polynomial activation on the Tiny-ImageNet dataset. In Corollary 4.3 and Corollary 4.5 it was pointed out that the scaling of the polynomial is important to keep the Frobenius norm of the polynomial attention matrix from becoming too large. The results of those corollaries suggested that a scale to use is $\frac{1}{\sqrt{N}}$, $N$ being the sequence length, which in this case is $\frac{1}{\sqrt{256}}=\frac{1}{16}$. We therefore decided to to compare the three activations softmax, $x^{3}$ and $\frac{1}{16}x^{3}$. Each was trained for 200 epochs using the AdamW optimizer. To begin with we computed the Frobenius norm of the attention matrix throughout training averaged over all the heads in layers 2, 7 and 12 of the ViT-Tiny architecture when trained on the Tiny-ImageNet dataset. Figure 2 plots the results. We see from the figure that the Frobenius norm of of the $x^{3}$ archicture is much larger than softmax but scaling it by $\frac{1}{16}$ brings it down to softmax levels. Similarly, Figure 3 shows the Jacobian’s Frobenius norm, where scaling also brings the norms closer to softmax, ensuring more stable gradients. Further plots for other layers are in Section A.2.3. Table 1 presents the final Top-1% accuracy achieved by each activation function. Notably, $\frac{x^{3}}{16}$ delivers the best performance. In contrast, the unscaled $x^{3}$ activation yields significantly lower Top-1% accuracy, underscoring the importance of incorporating an appropriate scaling factor.
@@ 6 Experiments
Building on the theoretical foundations from Section 4, we focus on the cubic polynomial $x^{3}$ as a test case. Notably, $x^{3}$ introduces both positive and negative values, does not normalize rows to sum to 1, and generally produces dense attention matrices—violating all three traditional softmax conditions.
@@
    Fixed scale: Following Section 4, we scale $x^{3}$ by the inverse square root of the sequence length (which is fixed throughout training), as theory suggests this maintains optimization stability.
@@
    Dynamic scale: Recognizing that the assumption of i.i.d. normal-distributed $Q$, $K$, $V$, used in Section 4, holds primarily at initialization, we also explore a learnable scale. This dynamic scale is initialized as above but optimized during training.
@@ 6.1 Image Classification
We trained all models on the ImageNet-1k dataset from scratch and report Top-1 accuracy on the validation set. We examined our approach along with the following four transformer architectures to show its generalization, ViT-B [10], DeiT [25], Swin Transformer [17], XCiT [3]. Each transformer was trained following the approach in each paper. The results are shown in Table 2. As can be seen from the table the $x^{3}$ activation with a dynamic scale performed the best and the one with a fixed scale performed comparable to softmax. On the other hand $x^{3}$ (with no scale) underperformed on all ViTs showing how important the theory on scaling as developed in Section 4 is. For a discussion on other polynomials and linear attention see Section A.2.4.
@@
Softmax attention traditionally satisfies three key properties: positivity, row normalization, and sparsity. In contrast, the polynomial activation $x^{3}$ (with or without positive scaling) takes positive values for $x>0$ and negative values for $x<0$, thus violating these softmax constraints. To better understand how this affects attention patterns, we analyzed the self-attention matrices of ViT-B models trained with the $x^{3}$ + dynamic scale activation and compared them to those using softmax. We visualized heatmaps of the attention matrices after convergence, focusing on two representative layers and heads, and averaged over a fixed batch of 128 samples. Figure 4 shows results for layer 2, head 8, where the $x^{3}$ + dynamic scale activation produces attention scores with both positive and negative values, unlike softmax. Similarly, Figure 5 illustrates distinct patterns for layer 12, head 6, highlighting how the two activations differ in learned attention distributions.
@@
Even with these differences, the $x^{3}$ + dynamic scale activation still learns meaningful attention patterns. For instance, in Figure 4 (left), we observe 14 bands about the diagonal, indicating that the head has learned to attend to patches in the same image row as the query patch, sufficient for effective image classification. This suggests that the sign of attention values (positive or negative) is not inherently critical for attention allocation. Further, Figure 5 (left) reveals vertical lines, showing attention that depends solely on key positions, independent of the query position. Smaller key indices receive higher attention weights, focusing model capacity where it matters most for classification tasks. These findings demonstrate that, despite deviating from softmax properties, the $x^{3}$ + dynamic scale activation enables the model to discover effective attention patterns. We noticed similar attention patterns for the $x^{3}$ + fixed scale activation.
@@ 6.2 Object Detection and Instance Segmentation
In this section, in order to examine the transfer learning ability of our models, we demonstrate our approach on object detection and segmentation tasks by fine-tuning a ImageNet-pretrained XCiT model on them. Our experiments are conducted on COCO 2017 [16], which has 118K training images and 5K validation images with 80 categories. We integrate the XCiT architecture as the backbone in the Mask R-CNN [13] detector with a Feature Pyramid Network (FPN). Due to XCiT’s inherently columnar design, we adapt it for FPN compatibility by extracting features from various layers for XCiT-S12. These features have a consistent stride of 16. The feature resolutions are then adjusted to strides of [4, 8, 16, 32]. This downsampling is accomplished through max pooling, while upsampling is achieved using a single transposed convolution layer. We conducted experiments on XCiT-S12 models using 16×16 patches with the activations $x^{3}$ + fixed scale, $x^{3}$ + dynamic scale, with scale initialized to $\frac{1}{14}$, and softmax. We found when we trained with $x^{3}$ the model would not converge thereby suggesting that a scaling is extremely important as predicted by the theory in Section 4.2. The training hyperparameters are given in Section A.2.2. Table 3 shows the results of the experiment. From the table we see that the activation $x^{3}$ + dynamic scale obtains better performance to softmax with $x^{3}$ + fixed scale obtaining comparable performance.
@@ 6.3 Nyströmformer on LRA benchmark
On each dataset we compared a Nyströmformer with softmax, $x^{3}$ + dynamic scale, and $x^{3}$. For the $x^{3}$ + dynamic scale activation on each dataset we used a different scale as initialization based on the sequence length used in the original Nyströmformer model. For ListOps the original model used a sequence length of $512$ and so the scale we used as initialization was $1/\sqrt{512}$, on Text Classification the input sequence used is $1024$ and so the scale we used at initialization was $1/\sqrt{1024}$, on Retrieval the input sequence used is $2048$ and the initialized scale we used was $1/\sqrt{2048}$, on Image Classification the input sequence is $4096$ and the initialized scale we used was $1/\sqrt{4096}$, and on Pathfinder the input sequence used is $8192$ and the initialized scale we used was $1/\sqrt{8192}$. We followed the training regime of the original paper [30]. The results are shown in Table 4. As can be seen from the table the $x^{3}$ + dynamic scale activation results in better performance compared to softmax with $x^{3}$ + fixed scale performing comparably. However $x^{3}$ (with no scale) underperforms on all tasks.
@@ 7 Physics Informed Transformers
Recently, dynamical data has become essential for modeling time-dependent systems. [31] introduced transformer architectures for solving partial differential equations (PDEs), demonstrating superior performance across various PDE tasks. These models effectively capture interactions between spatial and temporal components inherent in dynamical systems. In Section A.2.5, we extend our polynomial attention framework to Physics-Informed Transformers, showing that polynomial activations can outperform softmax in this setting.

### v3

@@ 3 Preliminaries and Notation
The factor $\frac{1}{\sqrt{d}}$, as explained in [26], is a scaling to prevent the gradients of softmax from being too small. For the theoretical analysis in this paper we will only use the dot-product similarity $qk^{T}$ and call the $N\times N$ matrix $softmax(qk^{T})$ the softmax self-attention matrix. In the experiments, Section˜6, we will empirically validate our theoretical framework on more general softmax attention blocks used in state of the art transformers such as DeiT [25], Swin Transformer [17] and XciT [30].
@@
Given a matrix $A\in\mathbb{R}^{n\times m}$, we denote its Frobenius norm by $A||_{F}$. Additionally, we use the notation $\mathbb{E}$ to represent the expectation of a random variable, where the specific random variable being considered will be clear from the context.
@@ 4.1 Implicit regulatization of Softmax
[…]
The key implication of theorem 4.1 is that during the training of a transformer with softmax self-attention, the Frobenius norm of each softmax self-attention matrix remains bounded by a value that grows as $\mathcal{O}(\sqrt{N})$. This ensures that backpropagation through the weights of the self-attention matrix does not lead to excessively large gradients. The proof hinges on the fact that the row normalization inherent in softmax effectively controls the Frobenius norm. For a detailed proof see Section˜A.1.1.
@@ 4.2 Polynomial activations for self-attention
In Section˜4.1, we demonstrated that softmax implicitly regularizes the Frobenius norm of the self-attention matrix. Building on this, we now show that by scaling specific polynomial activations, a similar regularization effect on the Frobenius norm can be achieved in expectation, closely replicating the impact of softmax.
@@
[…]
@@
[…]
Corollary˜4.3 establishes that activations of the form $\phi(x):=\frac{1}{\sqrt{N}}x^{p}$ provide a level of regularization, in expectation, similar to that of softmax when applied to the self-attention matrix. The proof of theorem 4.2 can be found in appendix A.1.2. The next property we want to prove is one similar to the gradient bound obtained in Theorem˜4.1. Since the self-attention matrix has parameters given by the queries $Q$ and keys $K$ [26], this implies that during the training of a transformer the $Q$ and $K$ matrices are the only aspects of the self-attention matrix that get updated. Therefore, we compute a derivative bound with respect to the $Q$ and $K$ derivatives.
@@
[…]
@@
[…]
@@
Corollary˜4.3 and Corollary˜4.5 suggest that polynomial activations of the form $\phi(x)=\frac{1}{\sqrt{N}}x^{p}$, with $p>0$, can achieve performance comparable to softmax when applied to self-attention matrices. We point out that both corollaries rested on the assumption that $X$, $Q$ and $K$ are i.i.d random variables. In general, this is only true at initialization when training a transformer [2]. Although this is a limitation in the theory the experiments in Section˜6 show that this insight can be used effectively to develop new attention blocks that perform comparable to softmax yet violate the three conditions of positivity, normalized rows summing to $1$ and sparsity showing that attention blocks do not need to be modeled as a probability distribution.
@@ 5 Testing the theory
In this section we test the theory developed in Section˜4.2 on small vision transformers. We will consider the activation $\phi(x)=\frac{1}{k}x^{3}$, where $k>0$ is a fixed scale, as this activation clearly violates the three key conditions of softmax based attention; positivity, normalization and sparsity. We found that polynomials $\frac{1}{k}x^{p}$ for $p>3$ did not perform well during training as they witnessed a gradient vanishing problem due to the fact that the function $x^{p}$ for $p$ large have very small values around 0.
@@
The first experiment we conducted was to test how the Top-$1\%$ accuracy changes for a ViT-Tiny vision transformer [22], trained on Tiny-Imagenet dataset [21], as we change the sequence length $N$ of the input and the scale predicted in corollaries 4.3 and 4.5 when using the activation $\frac{1}{k}x^{3}$. The standard ViT-Tiny model comprises $12$ transformer layers, each equipped with $3$ attention heads, with each head having dimension $64$. We considered four different input sequence lengths $N$ of sizes $256$, $64$, $16$ and $8$. For each such sequence length, we ran a ViT-Tiny architecture with the activation $\frac{1}{k}x^{3}$ where $k$ ranged from roughly $10^{-3}$ to $10^{3}$. According to the theory developed in section 4, the Frobenius norm of $\frac{1}{\sqrt{N}}x^{3}$ scales according to $\mathcal{O}(\sqrt{N})$. Thus the best accuracy should occur when $k=\mathcal{O}(\sqrt{N})$ and should degrade for other values due to training instability. Figure˜1 shows the results of the experiment, we note that the xaxisx-axis plots the values of $k$. We see that as the sequence length increases the factor of $k$ needs to increase so that the activation $\frac{1}{k}x^{3}$ performs well on the ViT-Tiny architecture as predicted by the theory developed in corollaries 4.3 and 4.5.
@@
In a second experiment we decided to compare the performance of the original ViT-Tiny architecture, that uses an input sequence length of $256$, with a softmax activation and a polynomial activation on the Tiny-ImageNet dataset. In Corollary˜4.3 and Corollary˜4.5 it was pointed out that the scaling of the polynomial is important to keep the Frobenius norm of the polynomial attention matrix from becoming too large. The results of those corollaries suggested that a scale to use is $\frac{1}{\sqrt{N}}$, $N$ being the sequence length, which in this case is $\frac{1}{\sqrt{256}}=\frac{1}{16}$. We therefore decided to to compare the three activations softmax, $x^{3}$ and $\frac{1}{16}x^{3}$. Each was trained for 200 epochs using the AdamW optimizer. To begin with we computed the Frobenius norm of the attention matrix throughout training averaged over all the heads in layers 2, 7 and 12 of the ViT-Tiny architecture when trained on the Tiny-ImageNet dataset. Figure˜2 plots the results. We see from the figure that the Frobenius norm of of the $x^{3}$ archicture is much larger than softmax but scaling it by $\frac{1}{16}$ brings it down to softmax levels. Similarly, Figure˜3 shows the Jacobian’s Frobenius norm, where scaling also brings the norms closer to softmax, ensuring more stable gradients. Further plots for other layers are in Section˜A.2.3. Table˜1 presents the final Top-1% accuracy achieved by each activation function. Notably, $\frac{x^{3}}{16}$ delivers the best performance. In contrast, the unscaled $x^{3}$ activation yields significantly lower Top-1% accuracy, underscoring the importance of incorporating an appropriate scaling factor.
@@ 6 Experiments
Building on the theoretical foundations from Section˜4, we focus on the cubic polynomial $x^{3}$ as a test case. Notably, $x^{3}$ introduces both positive and negative values, does not normalize rows to sum to 1, and generally produces dense attention matrices—violating all three traditional softmax conditions.
@@
    Fixed scale: Following Section˜4, we scale $x^{3}$ by the inverse square root of the sequence length (which is fixed throughout training), as theory suggests this maintains optimization stability.
@@
    Dynamic scale: Recognizing that the assumption of i.i.d. normal-distributed $Q$, $K$, $V$, used in Section˜4, holds primarily at initialization, we also explore a learnable scale. This dynamic scale is initialized as above but optimized during training.
@@ 6.1 Image Classification
We trained all models on the ImageNet-1k dataset from scratch and report Top-1 accuracy on the validation set. We examined our approach along with the following four transformer architectures to show its generalization, ViT-B [10], DeiT [25], Swin Transformer [17], XCiT [3]. Each transformer was trained following the approach in each paper. The results are shown in Table˜2. As can be seen from the table the $x^{3}$ activation with a dynamic scale performed the best and the one with a fixed scale performed comparable to softmax. On the other hand $x^{3}$ (with no scale) underperformed on all ViTs showing how important the theory on scaling as developed in Section˜4 is. For a discussion on other polynomials and linear attention see Section˜A.2.4.
@@
Softmax attention traditionally satisfies three key properties: positivity, row normalization, and sparsity. In contrast, the polynomial activation $x^{3}$ (with or without positive scaling) takes positive values for $x>0$ and negative values for $x<0$, thus violating these softmax constraints. To better understand how this affects attention patterns, we analyzed the self-attention matrices of ViT-B models trained with the $x^{3}$ + dynamic scale activation and compared them to those using softmax. We visualized heatmaps of the attention matrices after convergence, focusing on two representative layers and heads, and averaged over a fixed batch of 128 samples. Figure˜4 shows results for layer 2, head 8, where the $x^{3}$ + dynamic scale activation produces attention scores with both positive and negative values, unlike softmax. Similarly, Figure˜5 illustrates distinct patterns for layer 12, head 6, highlighting how the two activations differ in learned attention distributions.
@@
Even with these differences, the $x^{3}$ + dynamic scale activation still learns meaningful attention patterns. For instance, in Figure˜4 (left), we observe 14 bands about the diagonal, indicating that the head has learned to attend to patches in the same image row as the query patch, sufficient for effective image classification. This suggests that the sign of attention values (positive or negative) is not inherently critical for attention allocation. Further, Figure˜5 (left) reveals vertical lines, showing attention that depends solely on key positions, independent of the query position. Smaller key indices receive higher attention weights, focusing model capacity where it matters most for classification tasks. These findings demonstrate that, despite deviating from softmax properties, the $x^{3}$ + dynamic scale activation enables the model to discover effective attention patterns. We noticed similar attention patterns for the $x^{3}$ + fixed scale activation.
@@ 6.2 Object Detection and Instance Segmentation
In this section, in order to examine the transfer learning ability of our models, we demonstrate our approach on object detection and segmentation tasks by fine-tuning a ImageNet-pretrained XCiT model on them. Our experiments are conducted on COCO 2017 [16], which has 118K training images and 5K validation images with 80 categories. We integrate the XCiT architecture as the backbone in the Mask R-CNN [13] detector with a Feature Pyramid Network (FPN). Due to XCiT’s inherently columnar design, we adapt it for FPN compatibility by extracting features from various layers for XCiT-S12. These features have a consistent stride of 16. The feature resolutions are then adjusted to strides of [4, 8, 16, 32]. This downsampling is accomplished through max pooling, while upsampling is achieved using a single transposed convolution layer. We conducted experiments on XCiT-S12 models using 16×16 patches with the activations $x^{3}$ + fixed scale, $x^{3}$ + dynamic scale, with scale initialized to $\frac{1}{14}$, and softmax. We found when we trained with $x^{3}$ the model would not converge thereby suggesting that a scaling is extremely important as predicted by the theory in Section˜4.2. The training hyperparameters are given in Section˜A.2.2. Table˜3 shows the results of the experiment. From the table we see that the activation $x^{3}$ + dynamic scale obtains better performance to softmax with $x^{3}$ + fixed scale obtaining comparable performance.
@@ 6.3 Nyströmformer on LRA benchmark
On each dataset we compared a Nyströmformer with softmax, $x^{3}$ + dynamic scale, and $x^{3}$. For the $x^{3}$ + dynamic scale activation on each dataset we used a different scale as initialization based on the sequence length used in the original Nyströmformer model. For ListOps the original model used a sequence length of $512$ and so the scale we used as initialization was $1/\sqrt{512}$, on Text Classification the input sequence used is $1024$ and so the scale we used at initialization was $1/\sqrt{1024}$, on Retrieval the input sequence used is $2048$ and the initialized scale we used was $1/\sqrt{2048}$, on Image Classification the input sequence is $4096$ and the initialized scale we used was $1/\sqrt{4096}$, and on Pathfinder the input sequence used is $8192$ and the initialized scale we used was $1/\sqrt{8192}$. We followed the training regime of the original paper [30]. The results are shown in Table˜4. As can be seen from the table the $x^{3}$ + dynamic scale activation results in better performance compared to softmax with $x^{3}$ + fixed scale performing comparably. However $x^{3}$ (with no scale) underperforms on all tasks.
@@ 7 Physics Informed Transformers
Recently, dynamical data has become essential for modeling time-dependent systems. [31] introduced transformer architectures for solving partial differential equations (PDEs), demonstrating superior performance across various PDE tasks. These models effectively capture interactions between spatial and temporal components inherent in dynamical systems. In Section˜A.2.5, we extend our polynomial attention framework to Physics-Informed Transformers, showing that polynomial activations can outperform softmax in this setting.